OpenStudy (anonymous):

Simplify completely 25 w to the sixth power over 10 w to the third power divided by 30 w squared over 5 w

OpenStudy (debbieg):

Just reduce those common factors! and remember: \[\frac{ x^n }{ x^m }=x^{n-m}\]

OpenStudy (anonymous):

so it will be 15W

OpenStudy (anonymous):

15W^4

OpenStudy (jdoe0001):

\(\bf \Large \cfrac{ \frac{25w^6}{10w^3} }{ \frac{30w^2}{5w} } \ \ ?\)

OpenStudy (debbieg):

hmm... don't think so.... \[\frac{ 25w^6 }{10w^3 }\times\frac{ 5w }{30w^2 }=\frac{ 5w^3 }{2 }\times\frac{ 1 }{6w }=\frac{ 5w }{12}\]

OpenStudy (anonymous):

This is correct! B. 5x^2/12